pgarner:

Use iteration to solve for a, which is between 10 and 100; log(𝑎)/𝑎∙(𝑎−8) = .000346537

sillybilly123:

even in it's most primitive sense, i think you need tech coz you're solving \(\log a = \alpha \cdot a ( a- 8)\) so iteration-wise, as you are asking it, coding something up should still be pretty simple. there are those ancient Euler, Newton algorithms that are popular, and maybe that is what you are getting at. indeed, who knows what you are really asking about?!?!

Vocaloid:

Closed due to answered. Please re-open a new question if you need additional assistance. @

Vocaloid:

@pgarner